OpenStudy (anonymous):

help in this one solve 3z+3t=b-1 for z

OpenStudy (anonymous):

isolate z 3z + 3t = b - 1 -3t both sides 3z = b - 1 - 3t divide by 3 both sides z = b/3 - 1/3 - t

OpenStudy (mathlegend):

Can you stop giving out answers?

OpenStudy (mathlegend):

You know this is against the code of conduct @nphuongsun93

OpenStudy (mathlegend):

You are supposed to guide people to the answer.

OpenStudy (anonymous):

okay, was kind of lazy

OpenStudy (anonymous):

help solve the equations S/R+1=T/R for R in terms of S and T

OpenStudy (anonymous):

@yanci First you have to multiply both sides by R to get rid of denominator R Then isolate R

OpenStudy (anonymous):

so im left with S+1=T

OpenStudy (anonymous):

@yanci no, the correct one is: S+L=T now u can find the L

OpenStudy (anonymous):

is a 1 not a L

OpenStudy (anonymous):

@yanci oopz sorry, so this should be S+R=T now solve for R ;)

OpenStudy (anonymous):

i subtract R

OpenStudy (anonymous):

@yanci this is simple, u should subtract S from both sides, => S + R - S = T - S => R =

OpenStudy (anonymous):

is R=T-S

OpenStudy (anonymous):

@yanci yepz:)